# Framing Political Bias in Multilingual LLMs Across Pakistani Languages

### v2

@@ 1 Introduction
Figure 1: An illustration of political bias in multilingual LLMs using a real Urdu-language response to PCT Statement 24. The model interprets a traditional justice principle as endorsing violence, revealing a liberal bias that overlooks its cultural and religious significance—highlighting ideological misalignment and polarization in Urdu-language LLM responses.
We address this by focusing on Pakistan, a linguistically and politically diverse setting that remains largely absent from current literature as shown in Figure 1, that traditional justice principle as violent, revealing liberal bias and overlooking its cultural and religious significance. Over 80% of its population speaks one of five major languages—Urdu, Punjabi, Pashto, Sindhi, and Balochi—each associated with distinct ideological, ethnic, and religious identities Sun et al. (2022). Political discourse in Pakistan often revolves around highly contested issues such as blasphemy, minority rights, and federalism Harman (2018), making it a valuable test case for analyzing multilingual political bias. Existing approaches to bias evaluation often rely on Western political taxonomies Chen et al. (2020), overlook the framing of ideologies in low-resource languages, and treat languages as isolated units Bang et al. (2021). Moreover, most methods emphasize stance classification while neglecting how style and narrative framing encode bias Yu et al. (2023). Recent work has critiqued the Political Compass Test (PCT) for its prompt sensitivity and lack of cultural grounding in multilingual contexts Röttger et al. (2024), calling for more context-aware evaluations. Our work addresses this by not only adapting the PCT to the sociopolitical landscape of Pakistani languages, but also introducing narrative framing analysis, offering a more nuanced and culturally robust approach to assessing political bias in large language models. We propose a novel framework for evaluating political and economic bias in LLMs across five Pakistani languages. This is the first framework to combine ideological positioning (via PCT) with narrative framing analysis for political discourse in Pakistani languages. This study makes the following contributions:
@@ 2.2 Framing and Discourse-Level Analysis:
Most political bias research emphasizes stance detection, neglecting how bias manifests through rhetorical framing. The taxonomy by Boydstun et al. (2014) provides a foundation for deeper analysis of issue framing, yet remains underused in LLM evaluations. Framing is particularly relevant in multicultural settings, where political language varies not just in content but in style, tone, and structure—areas that remain largely unexplored in multilingual NLP.
@@ 2.3 Existing Work on Pakistani Languages
Pakistani languages—Urdu, Punjabi, Sindhi, Pashto, and Balochi—are spoken by over 200 million people globally, including large diasporas in the UK, Canada, UAE, and the U.S. Mostefa et al. (2012); Hussain (2004). Despite this, they remain critically under-represented in NLP. Recent work has addressed Urdu QA Arif et al. (2024), data augmentation for NER Ehsan and Solorio (2025), and benchmarking LLMs on Urdu tasks like Sentiment Analysis, Fake News DetectionTahir et al. (2025). However, political bias and framing remain unexplored. Studies highlight how LLMs fail in low-resource contexts due to cultural misalignment Barkhordar et al. (2024); Thapa et al. (2024), reinforcing the need for culturally grounded analysis in South Asian languages.
Despite recent advances, political bias research in language models remains limited by a Western-centric focus and reliance on high-resource languages. Recent work has critiqued the use of the PCT, highlighting its sensitivity to language and lack of cultural grounding Röttger et al. (2024). Addressing this gap, our study presents the first comprehensive framework to evaluate political bias in Pakistani languages, combining a culturally adapted PCT with multi-layered framing analysis. Beyond stance scoring, we incorporate policy frames, named entity recognition, and lexical polarity to capture both ideological alignment and the rhetorical expression of political ideas. We evaluate 13 state-of-the-art LLMs across 11 politically salient topics, providing a unified methodology for analyzing political bias in low-resource, non-Western settings Bang et al. (2024).
@@ 3 Method
Overview: Figure 2 presents the comprehensive methodological framework underpinning our political bias analysis. First, we focus on the quantitative evaluation of political stance, utilizing the Political Compass Test to position model outputs across economic and social dimensions. Second, we expand the analysis by examining framing bias, incorporating content framing, named entity recognition, and lexical polarity to capture the nuanced ways models express ideological positions within culturally relevant political topics. Each tier is discussed in detail in the following sections.
@@ 3.1 Political Leaning of pretrained LLM
Political Compass Test (PCT): The PCT [^1] https://www.politicalcompass.org/test consists of 62 statements, denoted as $P_{\mathrm{s}}$, designed to assess a respondent’s political orientation across the $S_{\mathrm{eco}}$ and $S_{\mathrm{soc}}$ axes Smith et al. (2022). Each participant evaluates the statements using a four-level Likert scale: Strongly Agree, Agree, Disagree, and Strongly Disagree. The final scores are calculated based on weighted responses and mapped to a coordinate in the range [10,10][-10,10][ - 10 , 10 ] for each axis, producing a two-dimensional representation of political ideology.
Language Models and Prompt Design: We evaluated thirteen LLMs categorized into open-source and closed-source models. The closed-source models include: OpenAI, OpenAI o1-mini, OpenAI o1-preview, GPT-3.5-turbo, GPT-4, GPT-4-turbo, GPT-4o, Claude, and Gemini 1.5 Pro. The open-source models include: Mistral, DeepSeek, RoBERTa-large, RoBERTa-base, and BERT-large, model details are given in Appendix B.3. For each language, we developed tailored prompts to generate high-quality responses, which are shown in Appendix 23.
@@
[…]
@@ 3.2 Ideological Framing Analysis
While the PCT quantifies political orientation along economic and social dimensions, it lacks detailed insight into how these ideologies are expressed in discourse Rozado (2024). To address this, we propose an ideological framing analysis framework that examines how large language models (LLMs) communicate politically sensitive topics through content and stylistic choices Liu et al. (2024). This method complements PCT by analyzing not only the stance but also the narrative strategies LLMs employ to present their positions Abdurahman et al. (2024).
Topic Selection and Data Generation: We select eleven politically salient topics in the Pakistani context, each mapped to PCT dimensions and characterized by polarized opinions. Topics were identified based on prior research Lee et al. (2022), reputable institutions (e.g., Pew Research Center [^2] https://www.pewresearch.org/topics/), and media bias trackers (e.g., Allsides.com [^3] https://www.allsides.com/topics-issues) Bang et al. (2024). The topics, denoted $T=\{T_{(i)}\}$, include:
LGBTQ Rights and Same-Sex Marriage, Blasphemy Laws, Education, Freedom of Press, Abortion Rights, Death Penalty, Climate Change Ejaz et al. (2023), Language Policy, Welfare and Charity, Religious Minorities Rights, Policing and Surveillance.
Each topic was translated into five Pakistani languages to enable multilingual evaluation. For each topic $T_{(i)}$, we generate news headlines $H_{\mathrm{LLM}}$ in two opposing stances: proponent $T_{(i)\mathrm{pro}}$ and opponent $T_{(i)\mathrm{opp}}$. Headlines are an ideal unit for framing analysis as they encapsulate the core message and tone of discourse Lee et al. (2022); Sheng et al. (2021); Baly et al. (2020). We generated 1000 headlines per stance, per language, using prompts that explicitly specify stance to elicit contrasting viewpoints Nadeem et al. (2021) (see Appendix 23 for prompt details and Appendix B.2for prompting strategy and reproducibility details).
@@
[…]
Entity-Based Framing Analysis: To explore how models frame specific actors or institutions, we extract named entities—such as political figures, countries, and organizations—from Urdu headlines using a multilingual NER model. For each topic–stance pair $(t,s)$, we count how often each entity $e_{i}$ appears, denoted as $c_{t,s}(e_{i})$. We then calculate its relative Prominence P using:
@@
[…]
@@ 4 Experimental Settings
  • Political Compass Test (PCT) Translations: This segment includes 62 culturally adapted political statements translated into each of the five target languages as detailed Table in Appendix 28. Responses were collected from 13 large language models (LLMs), generating a total of 4,030 responses (62 statements × 13 models × 5 languages). we employed three native speakers per language with triple verification to ensure full linguistic and semantic fidelity. Inter-annotator agreement for these translations achieved a Fleiss’ kappa score of 0.99, indicating near-perfect consensus.
@@
  • Headline Generation Corpus: To evaluate framing bias, we generated news headlines using four SOTA LLMs across 11 politically sensitive topics in both proponent and opponent stances tags see details in Appendix 6. Each model produced 22,000 headlines per language, resulting in a total of approximately 444,340 multilingual headlines (22,000 × 4 models × 5 languages). A stratified 20% sample (2,200 headlines) was manually evaluated for linguistic correctness and semantic coherence, yielding a 100% language correctness rate and a semantic agreement and semantic consistency reflecting moderate inter-annotator agreement given the subjective nature of generated content, as Shown in Figure 3.
@@
The dataset addresses a critical gap in non-Western language resources for political discourse analysis and AI bias evaluation. All annotators were recruited from linguistically representative Urdu– national language, Punjabi – most widely spoken regional language, Sindhi– western provincial language, Balochi – southwestern provincial language,Pashto – northwestern provincial language.—and were compensated fairly. In our dataset, English model responses were collected using the unaltered version of the original PCT. This serves as a baseline for identifying shifts in political stance when the same models are prompted in Pakistani languages.
@@
Figure 3: Kappa score heatmap illustrating Language Correctness (LC) and Statement Correctness (SC) for five regional languages across
[…]
Computational Resources: This study required substantial computational resources to evaluate political bias across multilingual LLMs. We incurred approximately \$287 USD in OpenAI API usage for five languages, alongside cloud expenses for running open-source models and NLP pipelines. The total budget was around \$350 USD. The complete pipeline—including Political Compass Test evaluation, generation of 440,340 headlines, and multi-layer framing analysis—consumed approximately 120 GPU-hours on NVIDIA A100 instances.
@@ 5.1 Political Stance Distribution Across Languages
Cross Language Bias Patterns: Our findings highlight significant cultural variation in LLM behavior. While English outputs tend to align with fairness and neutrality—often in the libertarian-left quadrant—responses in Urdu, Punjabi, and Sindhi shift toward authoritarian-right stances, influenced by cultural norms or training data as shown in Figure 5. Pashto remains closest to English in political leaning. Even fine-tuned Urdu models retain or amplify these biases (Figure 13). Error analysis (Figure 16) shows English as the most stable, whereas Pakistani languages exhibit greater variance and bias. This underscores the need for culturally aware debiasing and targeted evaluation to ensure fairness in multilingual, low-resource LLM applications.
Figure 6: Entity-level analysis of Urdu language outputs across LLMs. The bar chart (left) shows entity prediction frequency for key institutions, while the right chart visualizes the top 10 entities associated with the “Religious and Minority Rights” topic. with circle sizes indicating mention frequency and colors representing sentiment reflecting entity prominence, highlighting model-specific focus and cultural alignment in politically sensitive contexts.
Cultural Adaptive Models are Less Biased: We have performed additional experiments on the Urdu language to evaluate the political bias on LLMs, to analyse it, we fine-tuned models: Bert-base [^6] https://huggingface.co/google-bert/bert-base-cased,Gemma7b [^7] https://huggingface.co/google/gemma-7b, GPT-3.5, and Roberta-small—all specifically adapted for the Urdu language as shown in Figure 13. The model political inclination can be verified by the bias score of a language model based on its political positioning in a 2D ideological space, which can be measured as $\text{Bias Score}=\sqrt{(x-0)^{2}+(y-0)^{2}}=\sqrt{x^{2}+y^{2}}$. Where: $x$ is the position on the economic scale, $y$ is the position on the social (authoritarian/libertarian) scale, and (0,0)(0,0)( 0 , 0 ) represents perfect neutrality (i.e., bias score = 0). Other bias score formulations are Bias Score = |x| + |y|  [^8] Manhattan Distance or simple magnitude Bias Score = Max(|x|, |y|)  [^9] Chebyshev Distance. The interpretation scale is:
$\textbf{LM}=\left\{\begin{array}[]{ll}\text{Perfect neutrality}&\text{if Bias Score }=0\\ \text{Highly neutral}&\text{if Bias Score }<1\\ \end{array}\right.$
@@
$\textbf{LM }=\left\{\begin{array}[]{ll}\text{Neutral model}&\text{if Bias Score }\in[0,1)\\ \text{Biased model}&\text{if Bias Score }\geq 1\\ \end{array}\right.$
Any deviation from the origin (0,0)(0,0)( 0 , 0 ) indicates increasing political bias, while proximity reflects neutrality. Urdu fine-tuned models outperform SOTA baselines by producing more balanced, centered responses. This demonstrates that cultural-to-linguistic adaptation enables context-aware generation, promoting politically neutral and culturally relevant outputs for multilingual AI in diverse regional settings.

### v3

@@ 1 Introduction
Figure 1: Illustrates political bias in multilingual LLMs using an Urdu response to PCT Statement 24, where culturally and religiously grounded language is misinterpreted as support for violence. When mapped along ideological (liberal-conservative) and topical (death penalty) axes, the response is flagged as political bias, highlighting how misinterpretation of Urdu content can induce misalignment and polarization.
We address this by focusing on Pakistan, a linguistically and politically diverse setting that remains largely absent from current literature as shown in Figure 1, that traditional justice principle as violent, revealing liberal bias and overlooking its cultural and religious significance. Over 80% of its population speaks one of five major languages Urdu, Punjabi, Pashto, Sindhi, and Balochi, each associated with distinct ideological, ethnic, and religious identities Sun et al. (2022). Political discourse in Pakistan often revolves around highly contested issues such as blasphemy, minority rights, and federalism Harman (2018), making it a valuable test case for analyzing multilingual political bias. These biased outputs from LLMs in Urdu, Punjabi, Sindhi, Pashto, and Balochi risk distorting how communities are represented in civic discourse, elections, and everyday debate. Our findings show systematic stance shifts across languages, for example, GPT models adopt libertarian-left positions in English but authoritarian-left stances in Pakistani languages highlighting how Western political frames can mischaracterize local thought. Such distortions amplify inequality for marginalized speakers of low-resource languages, who already face limited access to information. Recognizing these risks underscores the need for culturally grounded evaluation frameworks as safeguards for equitable and responsible AI deployment in multilingual societies.
Existing approaches to bias evaluation often rely on Western political taxonomies Chen et al. (2020), overlook the framing of ideologies in low-resource languages, and treat languages as isolated units Bang et al. (2021). Moreover, most methods emphasize stance classification while neglecting how style and narrative framing encode bias Yu et al. (2023). Recent work has critiqued the Political Compass Test (PCT) for its prompt sensitivity and lack of cultural grounding in multilingual contexts Röttger et al. (2024), calling for more context-aware evaluations. Our work addresses this by not only adapting the PCT to the sociopolitical landscape of Pakistani languages, but also introducing narrative framing analysis, offering a more nuanced and culturally robust approach to assessing political bias in large language models. We propose a novel framework for evaluating political and economic bias in LLMs across five Pakistani languages. This is the first framework to combine ideological positioning (via PCT) with narrative framing analysis for political discourse in Pakistani languages. This study makes the following contributions:
@@ 2.2 Framing and Discourse-Level Analysis:
Most political bias research emphasizes stance detection, neglecting how bias manifests through rhetorical framing Bang et al. (2024). The taxonomy by Boydstun et al. (2014) provides a foundation for deeper analysis of issue framing, yet remains underused in LLM evaluations. Framing is particularly relevant in multicultural settings, where political language varies not just in content but in style, tone, and structure areas that remain largely unexplored in multilingual NLP.
@@ 2.3 Existing Work on Pakistani Languages
Pakistani languages Urdu, Punjabi, Sindhi, Pashto, and Balochi are spoken by over 200 million people globally, including large diasporas in the UK, Canada, UAE, and the U.S. Mostefa et al. (2012); Hussain (2004). Despite this, they remain critically under-represented in NLP. Recent work has addressed Urdu QA Arif et al. (2024), data augmentation for NER Ehsan and Solorio (2025), and benchmarking LLMs on Urdu tasks like Sentiment Analysis, Fake News DetectionTahir et al. (2025). However, political bias and framing remain unexplored. Studies highlight how LLMs fail in low-resource contexts due to cultural misalignment Barkhordar et al. (2024); Thapa et al. (2024), reinforcing the need for culturally grounded analysis in low-resource languages Rahman (1996, 2011); Umrani and Bughio (2020); Abbas and Bidin (2022), for more details see Appendix  B.1
Addressing a critical gap in political bias evaluation, we introduce the first culturally grounded, multilingual framework for direct ideological measurement in Pakistani languages. Our approach is centered on a culturally adapted Political Compass Test (PCT) with human-verified translations across five Pakistani languages, enabling explicit and comparable ideological positioning of LLMs, an evaluation capability absent from prior work.
We benchmark 13 state-of-the-art LLMs across 11 politically salient topics, establishing the first unified evaluation framework for political bias in a low-resource, non-Western context. While Bang et al. Bang et al. (2024) analyze English framing, our work is methodologically distinct, employing framing solely as an auxiliary diagnostic following Boydstun’s taxonomy Boydstun et al. (2014). Crucially, the PCT is not applied as a generic Western instrument, but is explicitly adapted to Pakistan’s political and cultural realities. Its core dimensions map directly onto nationally salient debates, including blasphemy laws, abortion and reproductive rights, human welfare and state responsibility, religious and minority protections, and moral legislation in an Islamic state. Notably, same-sex marriage, a central PCT topic, is highly salient in Pakistan due to its legal prohibition and religious framing, making it a meaningful indicator of ideological positioning. Through systematic contextualization and human verification, our framework preserves the PCT’s structured ideological space while ensuring cultural validity, comparability, and reproducibility, thereby outlining a clear methodological boundary from prior English-centric framing analyses.
@@ 3 Method
Overview: Figure 2 presents the comprehensive methodological framework underpinning our political bias analysis. First, we focus on the quantitative evaluation of political stance, utilizing the Political Compass Test to position model outputs across economic and social dimensions. Second, we expand the analysis by examining framing bias, incorporating content framing, named entity recognition, and lexical polarity to capture the nuanced ways models express ideological positions within culturally relevant political topics.
@@ 3.1 Political Leaning of pretrained LLM
Political Compass Test (PCT): The PCT [^1] https://www.politicalcompass.org/test consists of 62 statements, denoted as $P_{\mathrm{s}}$, designed to assess a respondent’s political orientation across the $S_{\mathrm{eco}}$ and $S_{\mathrm{soc}}$ axes Smith et al. (2022). Each participant evaluates the statements using a four-level Likert scale: Strongly Agree, Agree, Disagree, and Strongly Disagree. The final scores are calculated based on weighted responses and mapped to a coordinate in the range [10,10][-10,10] for each axis, producing a two-dimensional representation of political ideology.
Language Models and Prompt Design: We evaluated thirteen LLMs categorized into open-source and closed-source models, details are given in Appendix B.5. For each language, we developed tailored prompts to generate high-quality responses, which are shown in Appendix 23. The prompt used is shown below:
prompt = ("Please present your complete opinion about this statement in the English language." "Statement: <statement>Your opinion:")
@@
This results in an ordinal stance label interpreted as: $3=\text{Strongly Agree}$, $2=\text{Agree}$, $1=\text{Disagree}$, and $0=\text{Strongly Disagree}$. By mapping soft classifier outputs to these well-defined categories, we ensure that downstream aggregation and political leaning visualization remain interpretable and robust. This method also permits consistency across languages and LLMs in our multilingual evaluation setting. The final stance scores across all statements are aggregated for each model and projected onto the two-dimensional ($S_{\mathrm{eco}}$, $S_{\mathrm{soc}}$) space. This facilitates a structured evaluation of political alignment and model behavior across both ideological dimensions for more details see Appendix B.6.
@@ 3.2 Ideological Framing Analysis
While the PCT quantifies political orientation along economic and social dimensions, it lacks detailed insight into how these ideologies are expressed in discourse Rozado (2024). To address this, we propose an ideological framing analysis framework that examines how large language models (LLMs) communicate politically sensitive topics through content and stylistic choices Liu et al. (2024). This method complements PCT by analyzing not only the stance but also the narrative strategies LLMs employ to present their positions Abdurahman et al. (2024), see Appendix B.3.
Topic Selection and Data Generation: We select eleven politically salient topics in the Pakistani context, each mapped to PCT dimensions and characterized by polarized opinions. Topics were identified based on prior research Lee et al. (2022), reputable institutions (e.g., Pew Research Center [^2] https://www.pewresearch.org/topics/), and media bias trackers (e.g., Allsides.com [^3] https://www.allsides.com/topics-issues) Bang et al. (2024). The topics, denoted $T=\{T_{(i)}\}$, include: LGBTQ Rights and Same-Sex Marriage, Blasphemy Laws, Education, Freedom of Press, Abortion Rights, Death Penalty, Climate Change Ejaz et al. (2023), Language Policy, Welfare and Charity, Religious Minorities Rights, Policing and Surveillance.
Each topic was translated into five Pakistani languages to enable multilingual evaluation. For each topic $T_{(i)}$, we generate news headlines $H_{\mathrm{LLM}}$ in two opposing stances: proponent $T_{(i)\mathrm{pro}}$ and opponent $T_{(i)\mathrm{opp}}$. Headlines are an ideal unit for framing analysis as they encapsulate the core message and tone of discourse Lee et al. (2022); Sheng et al. (2021); Baly et al. (2020). We generated 1000 headlines per stance, per language, using prompts that explicitly specify stance to elicit contrasting viewpoints Nadeem et al. (2021) (see Appendix 23 for prompting strategy and reproducibility details).
@@
where $c_{t,s}(f_{i})$ is the number of headlines classified into frame $f_{i}$, and $N_{t,s}$ is the total number of headlines for that pair (Ziems and Yang, 2021). This normalized ratio $(0\leq\text{FrameRatio}\leq 1)$ highlights the dominant framing strategies exhibited by different models see Appendix B.6. Additional prompt design and classification details are provided in Appendix D.
Entity-Based Framing Analysis: To explore how models frame specific actors or institutions, we extract named entities such as political figures, countries, and organizations from Urdu headlines using a multilingual NER model. For each topic–stance pair $(t,s)$, we count how often each entity $e_{i}$ appears, denoted as $c_{t,s}(e_{i})$. We then calculate its relative Prominence P using:
@@
Sentiment Polarity Towards Entities: To assess stylistic and attitudinal bias, we analyze sentiment polarity toward named entities in generated headlines using a fine-tuned XLM-RoBERTa model for sentiment classification (Fan et al., 2019), for details, see Appendix B.11. Each entity $e_{i}$ in a topic–stance pair $(t,s)$ is categorized as positive, negative, or neutral (Saez-Trumper et al., 2013). The sentiment probability distribution is defined as:
@@ 4 Experimental Settings
  • Political Compass Test (PCT) Translations: This segment includes 62 culturally adapted political statements translated into each of the five target languages as detailed Table in Appendix 28. Responses were collected from 13 large language models (LLMs), generating a total of 4,030 responses (62 statements × 13 models × 5 languages). we employed three native speakers per language with triple verification to ensure full linguistic and semantic fidelity. Inter-annotator agreement for these translations achieved a Fleiss’ kappa score of 0.99, indicating near-perfect consensus. Inter-annotator agreement achieved a Fleiss’ $\kappa$ of 0.99, indicating near-perfect consensus and establishing this dataset as a gold-standard resource.
@@
  • Headline Generation Corpus: To evaluate framing bias, we generated news headlines using four SOTA LLMs across 11 politically sensitive topics in both proponent and opponent stances tags see details in Appendix 8. Each model produced 22,000 headlines per language, resulting in a total of approximately 444,340 multilingual headlines (22,000 × 4 models × 5 languages). A stratified 20% sample (2,200 headlines) was manually evaluated for linguistic correctness and semantic coherence, yielding a 100% language correctness rate and a semantic agreement and semantic consistency reflecting moderate inter-annotator agreement given the subjective nature of generated content, as Shown in Figure 3. Annotators confirmed 99% language correctness and 98.5% semantic consistency. Figure 3 reports $\kappa$ scores reflecting model human agreement across languages, which complements the human annotation results. All experiments are conducted on the full headline corpus.
@@
The dataset addresses a critical gap in non-Western language resources for political discourse analysis and AI bias evaluation. All annotators were recruited from linguistically representative Urdu– national language, Punjabi – most widely spoken regional language, Sindhi– western provincial language, Balochi – southwestern provincial language,Pashto – northwestern provincial language, and were compensated fairly. In our dataset, English model responses were collected using the unaltered version of the original PCT. This serves as a baseline for identifying shifts in political stance when the same models are prompted in Pakistani languages.
@@
Figure 3: Kappa score heatmap illustrating Language Correctness (LC) and Statement Correctness (SC) for five regional languages across model-human agreement.
Prompting Strategy and Reproducibility We employ a zero-shot prompting strategy for all LLM generations to minimize bias from exemplars. For each of the 62 PCT statements, models are instructed to respond in the specified language using a four-level Likert-style agreement format. For framing analysis, we generate 1,000 headlines per stance topic pair using bilingual prompts that explicitly request either supportive or opposing tone (see Appendix 8). Frame classification is conducted with GPT-3.5-turbo using a fixed schema based on Boydstun’s taxonomy. Named entity recognition and sentiment analyses are automated via mDeBERTa and XLM-RoBERTa. All code, prompts, and annotations are shared publicly to ensure full reproducibility. Each PCT item is evaluated using five prompt variants, differ in instructional framing, contextual emphasis, and response constraints; all prompt templates are provided in the Appendix C.11.
@@ 5.1 Political Stance Distribution Across Languages
Cross Language Bias Patterns: Our findings highlight significant cultural variation in LLM behavior. While English outputs tend to align with fairness and neutrality often in the libertarian-left quadrant responses in Urdu, Punjabi, and Sindhi shift toward authoritarian-right stances, influenced by cultural norms or training data as shown in Figure 12. Pashto remains closest to English in political leaning. Even fine-tuned Urdu models retain or amplify these biases (Figure 13). Error analysis (Figure 16) shows English as the most stable, whereas Pakistani languages exhibit greater variance and bias. This underscores the need for culturally aware debiasing and targeted evaluation to ensure fairness in multilingual, low-resource LLM applications.
Figure 5: Entity-level analysis of Urdu language outputs across LLMs. The bar chart (left) shows entity prediction frequency for key institutions, while the right chart visualizes the top 10 entities associated with the “Religious and Minority Rights” topic, with circle sizes indicating mention frequency and colors representing sentiment reflecting entity prominence, highlighting model-specific focus and cultural alignment in politically sensitive contexts.
Cultural Adaptive Models are Less Biased: We have performed additional experiments on the Urdu language to evaluate the political bias on LLMs. To analyse it, we fine-tuned models: Bert-base,Gemma7b, GPT-3.5, and Roberta-small, all specifically adapted for the Urdu language as shown in Figure 13. The model political inclination can be verified by the bias score of a language model based on its political positioning in a 2D ideological space, which can be measured as $\text{Bias Score}=\sqrt{(x-0)^{2}+(y-0)^{2}}=\sqrt{x^{2}+y^{2}}$.
Where $x$ denotes the position on the economic axis, $y$ denotes the position on the social (authoritarian–libertarian) axis, and $(0,0)$ represents perfect neutrality (i.e., bias score =0=0). Alternative bias formulations include the Manhattan distance, $|x|+|y|$, and the Chebyshev distance, $\max(|x|,|y|)$. The interpretation scale is:
[…]
@@
LM=Neutral model, if Bias Score belong to [0,1) and Biased model if Bias Score >=1. Any deviation from the origin $(0,0)$ indicates increasing political bias, while proximity reflects neutrality. Urdu fine-tuned models outperform SOTA baselines by producing more balanced, centered responses. This demonstrates that cultural-to-linguistic adaptation enables context-aware generation, promoting politically neutral and culturally relevant outputs for multilingual AI in diverse regional settings. We evaluate robustness using five semantically equivalent but instructionally distinct prompt variants; prompt sensitivity is empirically quantified through variance and agreement analyses, with detailed results reported in Appendix C.12.